OpenStudy (anonymous):

Give the solution of the equation in both exact form and decimal form (rounded to 3 decimal places!). 1=81tan(6x+5)-8 Exact Form:_____ Decimal Form:_____ Am I starting it correctly? 1=81tan(6x+5)-8 add 8 to both sides 9=81tan(6x+5) divide both sides by 81 1/9=tan(6x+5) from here, I'm not too sure how to continue :/ thanks!!

OpenStudy (amistre64):

so far so good yes exact is to undo the tan with the inverse tan function

OpenStudy (anonymous):

so 1/9=arctan(6x+5) ? :/ wait just wanna make sure haha arctan=tan^-1 right?

OpenStudy (amistre64):

tan' is a simple way to type it but yeah arctan= tan' 1/9 = tan(a) , tan' both sides tan'(1/9) = a ; the rest is simple algebra moves

OpenStudy (anonymous):

so now i plug in right? do i plug in for the values where theres an x?

OpenStudy (anonymous):

if so, it'd be like this? 1/9=tan(6x+5) 1/9=(6(arctan(1/9))+5) ? or is it something else?

OpenStudy (amistre64):

1/9=tan(6x+5) tan'(1/9) = tan'(tan(6x+5)) tan'(1/9) = 6x+5 , continue to solve for x

OpenStudy (anonymous):

ohh okay arctan(1/9)-5=6x (arctan(1/9)-5) / 6 = x ? would that be the exact form? :/

OpenStudy (amistre64):

that is the exact form yes

OpenStudy (anonymous):

x= -0.81489 so x is about -0.815 ?

OpenStudy (amistre64):

maybe, i dont have my calc with me :)

OpenStudy (amistre64):

looks fine tho

OpenStudy (anonymous):

haha okie :P i just put tan^-1 (1/9) - 5 and then divided it by 6 hehe :)

OpenStudy (anonymous):

okie yay thanks :)

OpenStudy (amistre64):

yep